RC meeting and Run Plan, Saturday March 11th
Return to the RC Meeting Calendar
- RC: Holly
- RC Daily report: https://logbooks.jlab.org/entry/4147575
Accelerator completed their RF work yesterday, and we resumed data-taking around 1:15pm. Some automated replay plots are missing from the logbook as we somehow filled up the disk space on the a-onl. This was quickly resolved when users checked and cleaned up their personal directories. Around 3pm, the south linac cryo plan went down. This event was recovered around mid-swing, but then there was a 1L03 issue which delayed the resumption of beam until about 9pm. Harp swipes were completed with an acceptable beam profile at the target (https://logbooks.jlab.org/entry/4147176). We completed beam checkout around 10:30pm and moved to production on pol-3He. A few DAQ issues arose overnight that were all resolved using the wiki instructions. Most of the beam overnight was around 33uA. Recently this morning they tried to give us 40uA again, but it looks pretty unstable so they are currently trying to find a more stable current (probably around 30-33uA).
Day summary: https://logbooks.jlab.org/entry/4147547 Owl summary: https://logbooks.jlab.org/entry/4147311 Swing summary: https://logbooks.jlab.org/entry/4147081 Day summary (yesterday): https://logbooks.jlab.org/entry/4147048
- While we request and are able to run at 45uA, accelerator has us running at 33uA for stability with Hall C online.
- Paul King noted that the asymmetry was leveling out around -200 ppm. https://logbooks.jlab.org/entry/4147493 He adjusted the PITA offset https://logbooks.jlab.org/entry/4147518
- Latest carbon hole image: https://logbooks.jlab.org/entry/4147220
- Running 45uA on pol-3He
- Production on 3He!
- GENII-3Stream config
- PS1=2, PS3=0, PS5=6
- Useful guide from Bill on rates to compare when bringing beam back: https://logbooks.jlab.org/entry/4144886
Data Collection and Polarization Status
- Currently 40%: https://logbooks.jlab.org/entry/4147448
- Data status at 4th-pass: 47.6 C (uncorrected)